OpenStudy (anonymous):

6.05] There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomore tickets totaled $400. How many sophomores bought tickets? 100 50 75 25

OpenStudy (anonymous):

first set up an equation for the problem

OpenStudy (anonymous):

do 2f + 3s =400 and f+s= 150

OpenStudy (anonymous):

Uh, I need help! :)

OpenStudy (anonymous):

do u understand the equation i set up

OpenStudy (anonymous):

Well, yes, I just don't know how to solve it. :(

OpenStudy (anonymous):

ok hang on im going to solve this then try to break this down step by step

OpenStudy (anonymous):

Ok, thx. :)

OpenStudy (anonymous):

ok so first you have to isolate f in order to find s so subtract f from the first equation

OpenStudy (anonymous):

what will that get you?

OpenStudy (anonymous):

Is the new equation 1f+2s=250

OpenStudy (anonymous):

yes but there are other ways to solve it try substitution, it will make more sense in the long run and will be much more efficient

OpenStudy (anonymous):

so your equation should be f=150-s

OpenStudy (anonymous):

so you can plug that back into your equation as 3(150-s) + 2s=400

OpenStudy (anonymous):

Is the answer 50?

OpenStudy (anonymous):

YES!

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

Thx, for your help!!!!!!!! OMG! Thx!

OpenStudy (anonymous):

no problem!